OpenStudy (anonymous):

The first line of paschen series for a certain hydrogen like species (ie an ion with only one electron) has a wavelenght 2085A identify the ion.?

OpenStudy (anonymous):

How do you calculate the wavelength of spectral lines in Hydrogen?

OpenStudy (anonymous):

Using rydberg formula

OpenStudy (anonymous):

Which is?

OpenStudy (anonymous):

1/lam=Rh*(1/n2-1/n'2)

OpenStudy (anonymous):

Rh=1/lam0

OpenStudy (anonymous):

\[ \frac{1}{\lambda} = Z^2\cdot(-13.6 eV) \cdot \left(\frac{1}{n^2} - \frac{1}{n'^2} \right) \]

OpenStudy (anonymous):

n = 3, n' = 4. Solve for Z.

OpenStudy (anonymous):

Yup..... i got it Z =3 Li ^2+ thxxxx